From: OPD Operations
Sent: Tue 10/25/2011 12:58 PM
To: OPD Operations; Israel, Jeffrey; Allison, Darren; Poirier, Michael; Santana, Deanna; Poulson, Edward; Parris, Kenneth; Garcia, Gilbert; Whent, Sean; Wiley, Kevin; Rachal, Anthony; Davis, Johnny; Lindsey, Drennon; Medeiros, Brian K.; Joyner, Ersie; Wong, Clifford W; Mestas, Fred; Lozares, Demetrio; Tull, Steven; Hamilton, Freddie; Williams, Sharon J; Shannon, Christopher; Outlaw, Danielle; DL – OPD – Sergeants; Jordan, Howard; Breshears, Eric
Situation Update #15 --- Operation FOP 25 OCT 2011 @ 1254 hours
2nd/Washington parking enforcement being followed by protestors, TNT responding. 14/bway 2 units responding here code 3, per dispatch there is a 943 here. Per Cpt Joyner enough units here to handle incident no need for more units. Patch Patrol 1/5 possibly being surrounded here. One I/C at 14/bway. Patrol units are being surrounded channel went code 33. Radio advised not to send any more patrol units here let units respond with helmets. Bravo 91 channel can be unpatched @1256hrs
Situation Update #19 --- Operation FOP 25 OCT 2011 @ 1428 hours
Per Chief Jordan protestors have found out the Mayors address, requesting units to be sent there to do a security check. Also giving AC Transit updates regarding reroutes.
Situation Update #21 --- Operation FOP 25 OCT 2011 @ 1603 hours
Another 40 people towards the library, total of 200 people at the library. Will provide update every 5 mins
Situation Update #23 --- Operation FOP 25 OCT 2011 @ 1614 hours
14/alice there is a march (80-100 people) headed eastbound in the middle of the street---also a group marching eastbound 14th, taking the street
Situation Update #25 --- Operation FOP 25 OCT 2011 @ 1620 hours
Crowd becoming big, starting to shut down traffic. 973 to post at all intersections north and south of 13th near the Library
Situation Update #27 --- Operation FOP 25 OCT 2011 @ 1627 hours
Everything is locked down on 14th st. No traffic coming from East Oakland along 14th st twds downtown. 15th st/Madison is where the hole is, this is where the traffic is. 17/Madison has stopped all south bound traffic.
Situation Update #28 --- Operation FOP 25 OCT 2011 @ 1700 hours
Crowd abt 600, 4-5 people masking up. 50-60 people on bikes. Snow Park is clear. Argus will be up in a minute or so. Bart is still up and running
Situation Update #30 --- Operation FOP 25 OCT 2011 @ 1736 hours
Lead of crowd w/b Webster. Santa Clara to stage inside of PAB. Argus advising not enough units at Webster. Will have police presence near the freeway entrance. Traffic needs to be stopped S/B 14th. Crowd is cooperative. 8 Motors ert to enforce @ 12/bway. 12/Bway overtaken.
Situation Update #31 --- Operation FOP 25 OCT 2011 @ 1757 hours
7/Bway blking traffic. All mobile field force relocating here. CHP sending 2 squads to ramps of freeway. Also broadcasted info regarding vigil in area of 18/mlk at the Church for 187 victim. Crowd headed westbound. Gilroy motors being sent to 8th/Bwy. Line held at 8th/bwy to make an arrest. Crowd headed w/b 7th street now. Per Argus officer is surrounded at 7th/Washington.
Situation Update #32 --- Operation FOP 25 OCT 2011 @ 1804 hours
7th/Washington was code 33 per Argus uncooperative protestors. Gas was deployed here also. Unit calling unlawful assembly. 7th/Clay appears to be peaceful. 8th/Clay protestors masking up. 7th/Washington unites will be forming a Scirmish Line. Hostiles are at 8th/Washington per Argus. Making 3rd announcement for unlawful assembly. Large group frm Washington/8th st
Situation Update #34 --- Operation FOP 25 OCT 2011 @ 1838 hours
Public works has pulled out of the plaza. Crowd moving northbound Bwy. Per Argus very little movement of crowd. 50-100 protestors starting to moving northbound Bwy. Group sitting down at 14/bwy. 2 Sheriff buses en route. Chp H30 taking over for Argus while they refuel. Per Argus 800-1000 protestors southbound 19/bwy. Tail of crowd at San Pablo. Holding now at 1700 Bwy taking over the street stopping traffic.
Situation Update #36 --- Operation FOP 25 OCT 2011 @ 1920 hours
Argus back on scene at Snow Park. Restablished barriers around the plaza. If anyone takes them down they are arrestable. Crowd is clearing Harrison approaching Webster. Crowd has grown 1200-1400 protestors at 20/bwy.
Situation Update #37 --- Operation FOP 25 OCT 2011 @ 1941 hours
Will allow protestors to peacefully protest. Not allowed to enter Frank Ogawa plaza if so they are arrestable. Crowd at 1500 Bwy holding. All units masking up making announcements at 14/Bwy given 5 mins to leave area apx 1500-1700 protestors here. Crown approaching barriers. Protestors have set up barriers, there is a bout a 7ft gap between their barriers and ours. Something being broken in the crowd, no visual. Distance now closed between barricades. Now throwing objects. Gas being deployed into crowd
Situation Update #38 --- Operation FOP 25 OCT 2011 @ 2016 hours
Numerous tweets that the Occupy Oakland General Assembly announced will meet everyday at 1800 hrs at 14th and Broadway (info via Sgt Dinh, T.) Male mid eastern on a red yamaha F-4 street bike w/ Camoflauge backpack handing out items to protestors in crowd. Trash cans being set on fire at 14/bwy. Protestor in crowd seen w/ a full triple CS de canister, suspect has wen back into the crowd. Holmgren advising bag w/ canister was recovered. Crowd now moving s/b San Pablo frm 20th. Argus down for fuel again, H30 (CHO) taking over.
Situation Update #39 --- Operation FOP 25 OCT 2011 @ 2058 hours
Fyi, arugs fueling up. Chp holding our spot. Upon our return, chp heading home, fog moving in and may be an issue. We will stay up as long as possible. Protestors throwing things Bwy ifo Rite Aid. Crowd size apx 600-700. Unlawful assembly announcement given. 16/Tel CHP vehicle had its rear window broken out and there is a weapon inside. 1200 blk of Bway protestors using dumpsters to make their own barricade. Unit took bottle at 14/Bway. Units have made 3.
Situation Update #40 --- Operation FOP 25 OCT 2011 @ hours
Argus coming back up. Checked weather, should not be a factor. Unit took another bottle 14/Bwy, protestors moving barricades. 1300 blk of Franklin apx 300 protestors. Protestors advising if we give them the park back we can go home. Protestors threw a bottle and hit their own people. Then threw several more. Making announcements now. Deploying gas
Situation Update #41 --- Operation FOP 25 OCT @ 2223hours
Crowd regathering at 15/Franklin. Crowd of 300-400 at 15/Franklin moving to 15/Bway (main crowd of apx 300). Small group 15/Franklin. Crowd moving northbound. 15/Bway deployed gas, were taking bottles here. Crowd moving westbound 17/bwy. Crowd has taken over intersections at 17/Bwy. apx 300 protestors. Snow park has apx 20-30 protestors here. Crowd moving southbound Bwy per Argus, still blking all lanes of traffic. 14/Bwy masking up again, crowd is reacting. Protestors are verbally challenging units on the line.
Situation Update #43 --- Operation FOP 25 OCT @ 2342hours
Unites were taking bottles, deployed gas again. Trash can on fire at 1624 Franklin. 15/Bwy apx 50-60 protestors spread out here. 14/Bwy 75-100, small group at 14/Franklin also apx 40-50. Units advd no fire at 17/Franklin. MW 6’0 ponytail, heavy padding w/backpack approaching the line, units advd to mask up.
— David Buuck : Oakland : 2011